Scores on a certain intelligence test for children between ages 13 and 15 years are approximately normally distributed with ?=106 and ?=21.

(a) What proportion of children aged 13 to 15 years old have scores on this test above 88 ? (Reminder: proportions are between 0 and 1 - don't put in percentages!)
Answer:

(b) Enter the score which marks the lowest 25 percent of the distribution.
Answer:

(c) Enter the score which marks the highest 5 percent of the distribution.
Answer:
Answer #1

Solution :

Given that ,

mean = \mu = 106

standard deviation = \sigma = 21

(a)

P(x > 88) = 1 - P(x < 88)

= 1 - P[(x - \mu ) / \sigma < (88 - 106) / 21]

= 1 - P(z <-0.8571 )

= 1 - 0.1957

= 0.8043

proportion = 0.8043

(b)

Using standard normal table ,

P(Z < z) =25%

P(Z < -0.67) = 0.25

z = -0.67

Using z-score formula,

x = z * \sigma + \mu

x = -0.67 * 21 + 106 = 91.93

Answer = 91.93

(c)

Using standard normal table ,

P(Z > z) = 5%

1 - P(Z < z) = 0.05

P(Z < z) = 1 - 0.05

P(Z < 1.65) = 0.95

z = 1.65

Using z-score formula,

x = z * \sigma + \mu

x = 1.65 * 21 + 106 = 140.65

Answer = 140.65